# Breaking the Gaussian Barrier: Residual-PAC Privacy for Automatic Privatization

### v1

@@ 6.1 A Stackelberg Game Model
Our SR-PAC algorithm recasts the optimal perturbation problem (8) as a two-level Stackelberg game between a Leader (choosing the perturbation rule $Q$) and Follower (choosing the decoder attempting to infer $X$ from $Y$). Let $\Gamma$ denote a rich family of noise distributions. Let $\Pi=\{\pi:\pi(\cdot|y)\in\Delta(\mathcal{X})\in\Delta(\mathcal{X}),y\in\mathcal{Y}\}$ denote a rich family of decoder distributions (e.g., all conditional density functions on $\mathcal{X}$ given $\mathcal{Y}$, or a parameterized neural-network family).
@@
$\inf_{\,Q\,\in\Gamma\,}\mathbb{E}_{X\sim P_{X},B\sim Q}\bigl{[}\mathcal{K}(B;\mathcal{M})\bigr{]},\text{ s.t. }\inf_{\pi\in\Pi}W(Q,\pi)\geq\hat{\beta}.$
@@
As stated in Proposition 4, a Stackelberg equilibrium perturbation rule solves (8).
@@
Algorithm 3 provides a Monte-Carlo-based approach to solve the Stackelberg equilibrium (9). By Monte Carlo sampling, this algorithm periodically trains the decoder to minimize reconstruction loss on perturbed data, enabling it to adapt to the current noise distribution. The perturbation rule is then optimized by balancing utility loss minimization against privacy constraints, using a penalty term that ensures the privacy cost remains close to the target budget.
@@ 6.2 Anisotropic Noise Perturbation
The Auto-PAC perturbs the mechanism using anisotropic Gaussian noise as much as needed in each direction of the output. This direction-dependent noise addition yields better privacy-utility tradeoffs than isotropic perturbation. Our SR-PAC also supports anisotropic noise generation under Assumption 1.
@@
    For any orthonormal direction $w\in\mathbb{R}^{d}$, $\mathcal{M}(X),w\right>$ is non-degenerate.
@@
    There exist orthonormal $u,v\in\mathbb{R}^{d}$ such that the marginal entropy gain per unit variance along $u$ exceeds that along $v$. That is, for any $\sigma^{2}>0$, $\frac{\partial}{\partial\sigma^{2}_{u}}\mathcal{H}(X|Z_{u})|_{\sigma^{2}}>\frac{\partial}{\partial\sigma^{2}_{v}}\mathcal{H}(X|Z_{v}|_{\sigma^{2}},$ where $Z_{w}=\mathcal{M}_{w}(X)+B_{w}$, with $A_{w}(X)=\left<A(X),w\right>$ for $A\in\{\mathcal{M},B\}$, $w\in\{u,v\}$.
@@
Proposition 5 demonstrates that SR-PAC strategically allocates noise exclusively to the most privacy-sensitive directions. By adjusting the covariance matrix so that high-leakage dimensions receive proportionally more noise while low-leakage dimensions receive less, the method ensures optimal noise utilization where each unit of perturbation yields maximum privacy protection. This targeted approach enables SR-PAC to achieve the desired privacy level with minimal total perturbation, thereby preserving task-relevant information with significantly reduced distortion.
@@ 6.3 Directional-Selectivity of SR-PAC
Let $Z\in\mathbb{R}^{d}$ be an output vector produced by a deterministic mechanism $\mathcal{M}(X)$; throughout we assume $\Sigma_{Z}\succ 0$ and finite differential entropy $\mathcal{H}(Z)$. For any application, let $S_{\mathrm{task}}\subseteq\mathbb{R}^{d}$ denote a practitioner‑chosen task‑critical sub‑space (the directions whose preservation matters most) and write $\Pi_{\mathrm{task}}$ for the orthogonal projector onto it.
Classification tasks. In what follows we illustrate the theory with multi‑class classification, where $Z$ is the logit vector, $\hat{y}=\arg\max_{i}Z_{i}$, and $S_{\mathtt{lab}}:=\operatorname{span}\{e_{\ell}-e_{j}:\,j\neq\ell\}$. Let $\Pi_{\mathtt{lab}}$ be the projector onto $S_{\mathtt{lab}}$. The analysis for a general $S_{\mathrm{task}}$ is identical after replacing $\mathtt{lab}$ by $\mathrm{task}$.
@@
For every unit vector $w$, let $g(w)\equiv\frac{1}{2}\mathtt{mmse}(\langle Z,w\rangle)$, where $\mathtt{mmse}(\langle Z,w\rangle)\equiv\mathbb{E}\Bigl{[}\bigl{\langle}Z,w\bigr{\rangle}\;-\;\mathbb{E}\bigl{[}\langle Z,w\rangle|Y\bigr{]}\Bigr{]}^{2}$ is the minimum mean‐squared error of estimating the scalar random variable $\langle Z,w\rangle$ from the noisy observation $Y=Z+B$.
@@
Fix any $0<\beta<\mathcal{H}(Z)$. The following holds.
@@
[…]
In Proposition 6, part (i) shows that SR-PAC always uses strictly less noise power than any Auto-PAC (regardless of how anisotropic the Auto-PAC noise covariance may be) because Auto-PAC treats $Z$ as Gaussian and thus overestimates the required variance when $Z$ is non-Gaussian. Part (ii) demonstrates that, under the natural ordering of directional sensitivities, SR-PAC allocates its noise budget exclusively in directions orthogonal to the label sub-space until a critical threshold $\beta_{\mathtt{lab}}$ is reached. In practice, this means SR-PAC perturbs only “harmless” dimensions first, preserving the predicted class and concentrating protection where it is most needed, thereby outperforming Auto-PAC in any scenario where certain directions leak more information than others.
@@ 6.4 Sensitivity to $\beta$
Sensitivity to the privacy parameter $\beta$ is crucial for predictable and accurate control of privacy-utility trade-off. Let $\mathtt{Priv}_{\beta}$ and $\mathtt{Util}_{\beta}$, respectively, denote the sensitivities of privacy and utility (for certain measures). If $\mathtt{Priv}_{\beta}=1$, then any infinitesimal increase $\Delta\beta$ in the privacy budget raises the true mutual information $\mathtt{MI}(X;Y)$ by exactly $\Delta\beta$. Thus, no part of the privacy budget is “wasted” or “over-consumed”. By contrast, if $\mathtt{Priv}_{\beta}<1$, then increasing $\beta$ may force additional noise without achieving the full allowed leakage; and if $\mathtt{Priv}_{\beta}>1$, even a small increase in $\beta$ could exceed the allowed privacy. Similarly, if $\mathtt{Util}{\beta}$ is high, then an infinitesimal increase $\Delta\beta$ in the privacy budget yields a large improvement in utility; if $\mathtt{Util}_{\beta}$ is low, the same increase yields a small improvement, indicating inefficient conversion of the privacy budget into utility gains.
@@
be the optimal noise-power curve attained by SR-PAC, and let $\mathtt{MI}_{\mathrm{SR}}(\beta)$ as the corresponding true mutual information attained by SR-PAC. Let $V_{\mathrm{PAC}}(\beta)\equiv\mathrm{tr}(\Sigma_{B_{\mathrm{PAC}}}(\beta))$, where $Q(\beta)=\mathcal{N}(0,\Sigma_{B_{\mathrm{PAC}}}(\beta))$ solves $\mathtt{LogDet}(\mathcal{M}(X),B_{\mathrm{PAC}})=\beta$. In addition, let $\mathtt{MI}_{\mathrm{PAC}}(\beta)\equiv\beta-\mathtt{Gap}_{\mathtt{d}}(Q(\beta)),$ where $\mathtt{Gap}_{\mathtt{d}}(Q)=\mathtt{D}_{\mathrm{KL}}(P_{\mathcal{M},B}\|\widetilde{Q}_{\mathcal{M}})$ with $B\sim Q$, and $\widetilde{Q}_{\mathcal{M}}$ given by (4). Define $\mathtt{Priv}^{\mathrm{SR}}_{\beta}\equiv\frac{d}{d\beta}\mathtt{MI}_{\mathrm{SR}}(\beta)$, $\mathtt{Priv}^{\mathrm{PAC}}_{\beta}\equiv\frac{d}{d\beta}\mathtt{MI}_{\mathrm{PAC}}(\beta)$, $\mathtt{Util}^{\mathrm{SR}}_{\beta}\equiv\frac{d}{d\beta}(-V_{\mathrm{SR}}(\beta))$, and $\mathtt{Util}^{\mathrm{PAC}}_{\beta}\equiv\frac{d}{d\beta}(-V_{\mathrm{PAC}}(\beta))$.
@@
Figure 1: Empirical comparisons of SR-PAC, Auto-PAC (Algorithm 1), and Efficient-PAC (Algorithm 2) on CIFAR-10, CIFAR-100, MNIST, and AG News as $\beta$ varies. ach column corresponds to one dataset; within each column, the three panels report (top) classification accuracy of the perturbed model versus the target budget $\beta$, (middle) the average noise power $\mathbb{E}[\lVert B\rVert_{2}^{2}]$ used by each method, and (bottom) the “target versus achieved” privacy budget for our SR-PAC.
@@ 6.5 Composition
In this section, we study the composition properties of Residual-PAC Privacy and SR-PAC, focusing on the conditional entropy formulation. Strong composition properties are essential for privacy definitions like differential privacy, enabling systems to quantify aggregated privacy risk across sequential, adaptive, and concurrent operations on related datasets. This allows modular design where individual components maintain local privacy-utility trade-offs while keeping global privacy risk quantifiable.
Consider $k$ mechanisms $\mathcal{M}_{1},\mathcal{M}_{2},\dots\mathcal{M}_{k}$, where each $\mathcal{M}_{i}(\cdot,\theta_{i}):\mathcal{X}\mapsto\mathcal{Y}_{i}$ with $\theta_{i}\in\Theta_{i}$ as the random seed. Let $\vec{\mathcal{Y}}=\prod^{k}_{i=1}\mathcal{Y}_{i}$ and let $\vec{\Theta}=\prod^{k}_{i=1}\Theta_{i}$. The composition $\overrightarrow{\mathcal{M}}(\cdot,\vec{\theta}):\mathcal{X}\mapsto\vec{\mathcal{Y}}$ is defined as $\overrightarrow{\mathcal{M}}(X,\vec{\theta})=\left(\mathcal{M}_{1}(X,\theta_{1}),\dots,\mathcal{M}_{k}(X,\theta_{k})\right).$ PAC Privacy composes gracefully. For independent mechanisms applied to the same dataset, mutual information bounds compose additively: if each $\mathcal{M}_{i}$ is PAC Private with bound $\beta_{i}$, then $\overrightarrow{\mathcal{M}}$ has bound $\sum^{k}_{i=1}\beta_{i}$.
Residual-PAC Privacy also enjoys additive composition. Suppose each mechanism $\mathcal{M}_{i}$ is Residual-PAC private with conditional entropy lower bound $\hat{\beta}_{i}$. By definition of mutual information, this implies that $\mathcal{M}_{i}$ is PAC private with privacy budget $\beta_{i}=\mathcal{H}(X)-\hat{\beta}_{i}$. Then, by Theorem 7 of [1], the composition $\overrightarrow{\mathcal{M}}(X,\vec{\theta})$ is PAC private with total mutual information upper bounded by $\sum_{i=1}^{k}(\mathcal{H}(X)-\hat{\beta}_{i})$. Equivalently, the composition $\overrightarrow{\mathcal{M}}(X,\vec{\theta})$ is Residual-PAC private with overall conditional entropy lower bounded by $\sum_{i=1}^{k}\hat{\beta}_{i}-(k-1)\mathcal{H}(X)$.
However, this additive composition property for mutual information yields conservative aggregated privacy bounds [1], and utility degradation compounds when each mechanism $\mathcal{M}_{i}$ uses conservative privacy budgets $\beta_{i}$. To address this limitation, we use an optimization-based approach within the SR-PAC framework to compute tighter conditional entropy bounds. Consider $k$ mechanisms $\mathcal{M}_{1},\mathcal{M}_{2},\dots\mathcal{M}_{k}$ privatized by $Q_{i}$ to satisfy Residual-PAC privacy with bounds $\hat{\beta}_{i}$. The Leader designs these privatizations $\{Q_{i}\}_{i=1}^{k}$, while the Follower finds the optimal decoder for the joint composition $\overrightarrow{\mathcal{M}}(X,\vec{\theta})=(\mathcal{M}_{1}(X),\dots,\mathcal{M}_{k}(X))$:
@@
7 Numerical Experiments
In this experiments, we use the followng four datasets: CIFAR-10 [40], CIFAR-100 [40], MNIST dataset [41], and AG News dataset [42].
@@ 7 Numerical Experiments
CIFAR-100 and Base Mechanism. CIFAR-100 contains 50,000 training and 10,000 testing color images (each $32\times 32\times 3$), equally divided among 100 fine‐grained classes (500 training and 100 testing images per class). Each image is converted to a $3\times 32\times 32$ tensor and normalized per channel to mean 0.5 and standard deviation 0.5 before being fed into the network. As the unperturbed mechanism, we use a deeper convolutional neural network with three convolutional “blocks.” Each block consists of two $3\times 3$ convolutions (with BatchNorm and ReLU after each), followed by a $2\times 2$ max‐pool, which sequentially maps inputs from $\times 32$ 3216×$\times 16$ 168×$\times 8$ 84×$\times 4$ 4, with channel widths increasing from 3 → 64 → 128 → 256. After flattening the resulting $256\times 4\times 4$ feature map into a 4096‐dimensional vector, a three‐layer MLP head (4096→512→256→100) with ReLU activations and 0.5 dropout between the first two fully connected layers produces a 100‐dimensional logit vector. During training, this network minimizes cross‐entropy loss over the CIFAR-100 classes; at inference, each normalized image is mapped to its 100-dimensional logits, and the predicted label is given by the $\arg\max$ of those logits. The unperturbed mechanism achieves 0.5913 accuracy.
@@
MNIST dataset and Base Mechanism. The MNIST dataset comprises 60,000 training and 10,000 test grayscale images of handwritten digits (0–9). Each image is $28\times 28$ pixels and is loaded as a $1\times 28\times 28$ tensor, then normalized to mean $0.1307$ and standard deviation $0.3081$ per channel before being fed into the network. As the unperturbed mechanism, we train a simple CNN consisting of two convolutional blocks—each block is Conv2d→BatchNorm→ReLU→MaxPool ($2\times 2$), with channel widths $1\to 32\to 64$—which produces a $64\times 7\times 7$ feature map. This feature map is flattened and passed through a two-layer fully connected head (128 units with ReLU+Dropout, then 10 output logits). At inference, each normalized $28\times 28$ image is mapped to a 10-dimensional logit vector, and the predicted label is given by the index of the largest logit. The unperturbed mechanism achieves 0.9913 accuracy.
@@
AG News dataset and Base Classifier. AG News comprises 120,000 training and 7,600 test articles equally divided among four classes (World, Sports, Business, Sci/Tech), i.e., 30,000 training and 1,900 test examples per class. Each example’s title and description are concatenated into one text string, then lowercased and split on whitespace (truncated or padded to 64 tokens). We build a 30,000‐word vocabulary from the training split and map each token to its index (with out‐of‐vocabulary tokens as 0). Those indices feed into an nn.EmbeddingBag layer (embedding size 300, mean‐pooling mode) to produce a fixed‐length 300‐dimensional document vector. That vector is passed through a two‐layer MLP head (300→256 with ReLU and 0.3 dropout, then 256→4), yielding a 4‐dimensional logit vector, and at inference the predicted label is the index of the largest logit. The unperturbed mechanism achieves 0.9729 accuracy.
@@
We evaluate SR-PAC against two baselines, Auto-PAC (Algorithm 1) and Efficient-PAC (Algorithm 2), on the above four datasets. For each dataset and its pretrained base mechanism $\mathcal{M}$, we plot (1) the test accuracy of the perturbed model as a function of $\beta$, (2) the average noise power $\mathbb{E}[\lVert B\rVert_{2}^{2}]$ required to achieve each $\beta$, and (3) SR-PAC’s ability to hit the target budget ^$\widehat{\mathcal{H}}(X)-\beta$ (where $\widehat{H}(X)$ is our plug-in entropy estimate using the non-parametric k-nearest neighbour approach). Figure 1 summarizes these comparisons.
Recall that $\beta$ is the user’s desired upper bound on the mutual information $\mathtt{MI}(X;M(X)+B)$. By construction, SR-PAC enforces the equivalent conditional-entropy constraint $\mathcal{H}\bigl{(}X\mid M(X)+B\bigr{)}\geq\widehat{\beta}=\mathcal{H}(X)-\beta$, but a direct per-$\beta$ comparison of accuracy would require knowing the true data entropy $\mathcal{H}(X)$, which is unavailable in practice. Fortunately, for any fixed mechanism $\mathcal{M}$ and dataset there is a single “intrinsic” mutual information $\mathtt{MI}_{o}=\mathtt{MI}(X;M(X)+B),$ and no independent noise $B$ can push $\mathtt{MI}(X;M(X)+B)$ above $\mathtt{MI}_{o}$. Equivalently, the only feasible budgets satisfy $0<\beta\leq\mathtt{MI}_{o}$. At the left endpoint $\beta=\mathtt{MI}_{o}$, the best solution is $B=0$, so all three methods coincide at the (near) noiseless accuracy. Although $\mathtt{MI}_{o}$ is unknown (and thus $\mathcal{H}(X)$ cannot be recovered), each algorithm still operates over the same interval $(0,\mathtt{MI}_{o}]$. In particular, every plotted $\beta$ is a valid privacy target for Auto-PAC, Efficient-PAC, and SR-PAC. Therefore, even without knowing $\mathtt{MI}_{o}$ (or $\mathcal{H}(X)$), we can fairly compare all three curves—accuracy vs. $\beta$ and noise magnitude vs. $\beta$—across the entire feasible range.
Accuracy vs. $\beta$ (a-d of Figure 1): As $\beta$ decreases (moving left along the horizontal axis), privacy increases. Thus, all three methods incur a drop in test accuracy. When $\beta$ is relatively large (near $\mathtt{MI}{o}$), all three algorithms yield almost the same accuracy close to the noiseless accuracy. As $\beta$ decreases, the SR-PAC curve always stays about the curves of Auto-PAC and Efficient-PAC—most dramatically on CIFAR-10 and CIFAR-100.
@@
Noise Magnitude vs. ($\beta$ (e-h) of Figure 1): As $\beta$ decreases, each algorithm must add more noise, so all three curves rise. In every dataset, SR-PAC uses the (approximately) smallest $\mathbb{E}[\|B\|^{2}_{2}]$ at each $\beta$, whose curves never exceed the blue and green curves. Auto-PAC and Efficient-PAC both overshoot. That is, they add strictly more noise than SR-PAC.

### v2

@@ 4.2 Stackelberg Automatic Residual-PAC Privatization
0: Privacy budget $\hat{\beta}$, decoder family $\Pi_{\phi}$, perturbation rule family $\Gamma_{\lambda}$, utility loss $\mathcal{K}(\cdot)$, learning rates $\eta_{\phi},\eta_{\lambda}$, penalty weight $\sigma$, iterations $T_{\lambda},T_{\phi}$, batch size $m$
@@
Our SR-PAC algorithm recasts the optimal perturbation problem (12) as a Stackelberg game between a Leader (who chooses the perturbation rule $Q$) and a Follower (who chooses the decoder attempting to infer $X$ from $Y$). Let $\Gamma$ denote a rich family of noise distributions. Let $\Pi=\{\pi:\pi(\cdot|y)\in\Delta(\mathcal{X}),y\in\mathcal{Y}\}$ denote a rich family of decoder distributions (e.g., all conditional density functions on $\mathcal{X}$ given $\mathcal{Y}$, or a parameterized neural network family).
@@
$\inf_{\,Q\,\in\Gamma\,}\mathbb{E}_{X\sim\mathcal{D},B\sim Q}\bigl{[}\mathcal{K}(B;\mathcal{M})\bigr{]},\text{ s.t. }\inf_{\pi\in\Pi}W(Q,\pi)\geq\hat{\beta}$.
@@
Proposition 5 shows that the Stackelberg equilibrium perturbation rule solves (12).
@@
Algorithm 2 provides a Monte-Carlo-based approach to solve the Stackelberg equilibrium (13). By Monte Carlo sampling, this algorithm periodically trains the decoder to minimize reconstruction loss on perturbed data, enabling it to adapt to the current noise distribution. The perturbation rule is then optimized by balancing utility loss minimization against privacy constraints, using a penalty term that ensures the privacy cost remains close to the target privacy budget.
5 Properties of SR-PAC Privatization
This section presents some important properties of SR-PAC.
@@ 5.1 Anisotropic Noise Perturbation
The Auto-PAC perturbs the mechanism using anisotropic Gaussian noise as much as needed in each direction of the output. This direction-dependent noise addition yields better privacy-utility tradeoffs than isotropic perturbation. SR-PAC also supports anisotropic perturbation under Assumption 1.
@@
    For any orthonormal direction $w\in\mathbb{R}^{d}$, $\left<\mathcal{M}(X),w\right$> is non-degenerate.
@@
    There exist orthonormal $u,v\in\mathbb{R}^{d}$ such that the marginal entropy gain per unit variance along $u$ exceeds that along $v$. That is, for any $\sigma^{2}>0$, $\frac{\partial}{\partial\sigma^{2}_{u}}\mathcal{H}(X|Z_{u})|_{\sigma^{2}}>\frac{\partial}{\partial\sigma^{2}_{v}}\mathcal{H}(X|Z_{v})|_{\sigma^{2}},$ where $Z_{w}=\mathcal{M}_{w}(X)+B_{w}$, with $A_{w}(X)=\left<A(X),w\right$> for $A\in\{\mathcal{M},B\}$, $w\in\{u,v\}$.
@@
Proposition 6 demonstrates that SR-PAC allocates noise exclusively to privacy-sensitive directions, with high-leakage dimensions receiving proportionally more noise than low-leakage dimensions. This targeted approach achieves desired privacy levels with minimal total perturbation, preserving task-relevant information with reduced distortion.
@@ 5.2 Directional-Selectivity of SR-PAC
Let $Z\in\mathbb{R}^{d}$ be an output vector produced by a deterministic mechanism $\mathcal{M}(X)$; throughout we assume $\Sigma_{Z}\succ 0$ and finite differential entropy $\mathcal{H}(Z)$. For any application, let $S_{\mathrm{task}}\subseteq\mathbb{R}^{d}$ denote a practitioner-chosen task-critical sub-space (the directions whose preservation matters most) and write $\Pi_{\mathrm{task}}$ for the orthogonal projector onto it.
Classification tasks. In what follows we illustrate the theory with multi-class classification, where $Z$ is the logit vector, $\hat{y}=\arg\max_{i}Z_{i}$, and $S_{\mathtt{lab}}:=\operatorname{span}\{e_{\ell}-e_{j}:\,j\neq\ell\}$, where $\mathtt{lab}$ means "label". Let $\Pi_{\mathtt{lab}}$ be the projector onto $S_{\mathtt{lab}}$. The analysis for a general $S_{\mathrm{task}}$ is identical after replacing $\mathtt{lab}$ by $\mathrm{task}$.
@@
For every unit vector $w$, let $g(w)\equiv\frac{1}{2}\mathtt{mmse}(\langle Z,w\rangle)$, where $\mathtt{mmse}(\langle Z,w\rangle)\equiv\mathbb{E}\Bigl{[}\bigl{\langle}Z,w\bigr{\rangle}\;-\;\mathbb{E}\bigl{[}\langle Z,w\rangle|Y\bigr{]}\Bigr{]}^{2}$ is the minimum mean-squared error of estimating the scalar random variable $\langle Z,w\rangle$ from the noisy observation $Y=Z+B$.
@@
Suppose $\mathcal{H}(Z)$ is finite. Fix any $0<\beta<\mathcal{H}(Z)$. The following holds.
@@
    Suppose $\sup_{v\in S_{\mathtt{lab}},\|v\|=1}g(v)<\inf_{w\perp S_{\mathtt{lab}},\|w\|=1}g(w)$. Let $\beta_{\mathtt{lab}}\equiv\frac{1}{2}\int_{w\bot S_{\mathtt{lab}}}g(w)d\sigma^{2}_{w}$ be the maximal MI reduction achievable with noise orthogonal to $S_{\mathtt{lab}}$. Then, for every $\beta\leq\beta_{\mathtt{lab}}$, we have $\Pi_{\mathtt{lab}}B^{*}=0\ \text{ a.s.},\;\arg\max_{i}(Z_{i}+B^{*}_{i})=\hat{y}\ \text{ a.s.}$
In Proposition 7, part (i) shows that SR-PAC always uses strictly less noise magnitude than any Auto-PAC (regardless of how anisotropic the Auto-PAC noise covariance may be) because Auto-PAC treats $Z$ as Gaussian and thus overestimates the required variance when $Z$ is non-Gaussian. Part (ii) demonstrates that, under the natural ordering of directional sensitivities, SR-PAC allocates its noise budget exclusively in directions orthogonal to the label sub-space until a critical threshold $\beta_{\mathtt{lab}}$ is reached. In practice, this means SR-PAC perturbs only "harmless" dimensions first, preserving the predicted class and concentrating protection where it is most needed, thereby outperforming Auto-PAC in any scenario where certain directions leak more information than others.
@@ 5.3 Sensitivity to $\beta$
Sensitivity to the privacy parameter $\beta$ is crucial for predictable and accurate control of privacy-utility trade-off. Let $\mathtt{Priv}_{\beta}$ and $\mathtt{Util}_{\beta}$, respectively, denote the sensitivities of privacy and utility (for certain measures). If $\mathtt{Priv}_{\beta}=1$, then any infinitesimal increase $\Delta\beta$ in the privacy budget raises the true mutual information $\mathtt{MI}(X;Y)$ by exactly $\Delta\beta$. Thus, no part of the privacy budget is “wasted" or “over-consumed". By contrast, if $\mathtt{Priv}_{\beta}<1$, then increasing $\beta$ may force additional noise without achieving the full allowed leakage; and if $\mathtt{Priv}_{\beta}>1$, even a small increase in $\beta$ could exceed the allowed privacy. Similarly, if $\mathtt{Util}{\beta}$ is high, then an infinitesimal increase $\Delta\beta$ in the privacy budget yields a large improvement in utility; if $\mathtt{Util}_{\beta}$ is low, the same increase yields a small improvement, indicating inefficient conversion of the privacy budget into utility gains.
@@
[…]
@@
Figure 1: Empirical comparisons of SR-PAC, Auto-PAC (Algorithm 1), and Efficient-PAC (Algorithm 3) on CIFAR-10, CIFAR-100, MNIST, and AG-News as $\beta$ varies. Each column corresponds to one dataset; within each column, the three panels report (top) classification accuracy of the perturbed model versus the target budget $\beta$, (middle) the average noise magnitude $\mathbb{E}[\lVert B\rVert_{2}^{2}]$ used by each method, and (bottom) the "target versus achieved" privacy budget (conditional entropy) for our SR-PAC.
@@ 5.4 Composition
[…]
Then, by Theorem 7 of Xiao et al. (2023), the composition $\overrightarrow{\mathcal{M}}(X,\vec{\theta})$ is PAC private with total mutual information upper bounded by $\sum_{i=1}^{k}(\mathcal{H}(X)-\hat{\beta}_{i})$. Equivalently, the composition $\overrightarrow{\mathcal{M}}(X,\vec{\theta})$ is R-PAC private with overall conditional entropy lower bounded by $\sum_{i=1}^{k}\hat{\beta}_{i}-(k-1)\mathcal{H}(X)$.
However, this additive composition property for mutual information yields conservative aggregated privacy bounds, and utility degradation compounds when each mechanism $\mathcal{M}_{i}$ uses conservative privacy budgets $\beta_{i}$. To address this limitation, we employ an optimization-based approach within the SR-PAC framework to compute tighter conditional entropy bounds. Consider $k$ mechanisms $\mathcal{M}_{1},\mathcal{M}_{2},\dots,\mathcal{M}_{k}$ privatized by distributions $Q_{i}$ to satisfy R-PAC privacy with bounds $\hat{\beta}_{i}$. The Leader designs these privatizations $\{Q_{i}\}_{i=1}^{k}$, while the Follower finds the optimal decoder for the joint composition $\overrightarrow{\mathcal{M}}(X,\vec{\theta})=(\mathcal{M}_{1}(X),\dots,\mathcal{M}_{k}(X))$:
@@
This game-theoretic formulation allows for tighter privacy-utility trade-offs in composed systems by optimizing the joint privatization strategy.
6 Experiments
(a) Iris
(b) Rice
Figure 2: Empirical comparisons of DP, Auto-PAC, Efficient-PAC, and SR-PAC on mean estimations, using Iris and Rice datasets, in terms of average noise magnitude $\mathbb{E}[\|B\|^{2}_{2}]$. All the numerical values are shown in Tables 5 and 6.
We conduct two sets of experiments to evaluate our approach. First, we compare SR-PAC against Auto-PAC and Efficient-PAC (Appendix A) using CIFAR-10 [29], CIFAR-100 [29], MNIST [31], and AG-News [48] datasets, with results presented in Section 6.1. We use (R-)PAC to refer to the family of SR-PAC, Auto-PAC, and Efficient-PAC. Second, we extend this comparison to include DP by equalizing optimal posterior success rates of membership inference (Appendix C.2) across all four methods (SR-PAC, Auto-PAC, Efficient-PAC, and DP), making their privacy budgets comparable. For this comparison, we use Iris [17] and Rice [9] datasets, with results shown in Section 6.2. All experiments focus on output perturbation. Appendix S provides more details of the experiments.
@@ 6 Experiments
CIFAR-100 and Base Classifier. CIFAR-100 contains 50,000 training and 10,000 testing color images (each $32\times 32\times 3$), equally divided among 100 fine-grained classes (500 training and 100 testing images per class). Each image is converted to a $3\times 32\times 32$ tensor and normalized per channel to mean 0.5 and standard deviation 0.5 before being fed into the network. The unperturbed classifier is a deep convolutional neural network with three convolutional “blocks.” Each block consists of two $3\times 3$ convolutions (with BatchNorm and ReLU after each), followed by a $2\times 2$ max-pool, which sequentially maps inputs from $32\times 32$$16\times 16$$8\times 8$$4\times 4$, with channel widths increasing from 3 → 64 → 128 → 256. After flattening the resulting $256\times 4\times 4$ feature map into a 4096-dimensional vector, a three-layer MLP head (4096→512→256→100) with ReLU activations and 0.5 dropout between the first two fully connected layers produces a 100-dimensional logit vector. During training, this network minimizes cross-entropy loss over the CIFAR-100 classes; at inference, each normalized image is mapped to its 100-dimensional logits, and the predicted label is given by the $\arg\max$ of those logits. The unperturbed classifier achieves $0.5914\pm 0.0090$ accuracy.
@@
MNIST dataset and Base Classifier. The MNIST dataset comprises 60,000 training and 10,000 test grayscale images of handwritten digits (0–9). Each image is $28\times 28$ pixels and is loaded as a $1\times 28\times 28$ tensor, then normalized to mean $0.1307$ and standard deviation $0.3081$ per channel before being fed into the network. The unperturbed classifier is a simple CNN consisting of two convolutional blocks—each block is Conv2d→BatchNorm→ReLU→MaxPool ($2\times 2$), with channel widths $1\to 32\to 64$—which produces a $64\times 7\times 7$ feature map. This feature map is flattened and passed through a two-layer fully connected head (128 units with ReLU+Dropout, then 10 output logits). At inference, each normalized $28\times 28$ image is mapped to a 10-dimensional logit vector, and the predicted label is given by the index of the largest logit. The unperturbed classifier achieves $0.9837$ accuracy.
@@
AG-News dataset and Base Classifier. AG-News comprises 120,000 training and 7,600 test articles equally divided among four classes (World, Sports, Business, Sci/Tech), i.e., 30,000 training and 1,900 test examples per class. Each example’s title and description are concatenated into one text string, then lowercased and split on whitespace (truncated or padded to 64 tokens). We build a 30,000-word vocabulary from the training split and map each token to its index (with out-of-vocabulary tokens as 0). Those indices feed into an nn.EmbeddingBag layer (embedding size 300, mean-pooling mode) to produce a fixed-length 300-dimensional document vector. That vector is passed through a two-layer MLP head (300→256 with ReLU and 0.3 dropout, then 256→4), yielding a 4-dimensional logit vector, and at inference the predicted label is the index of the largest logit. The unperturbed mechanism achieves $0.9705$ accuracy.
[…]
6.1 (R-)PAC Comparison
@@
For each dataset and its pretrained base classifier $\mathcal{M}$, we plot (1) the test accuracy of the perturbed model as a function of $\beta$, (2) the average noise magnitude $\mathbb{E}[\lVert B\rVert_{2}^{2}]$ required to achieve each $\beta$, and (3) SR-PAC’s ability to hit the target budget $\widehat{\mathcal{H}}(X)-\beta$ (where $\widehat{\mathcal{H}}(X)$ is our entropy estimate).
Accuracy vs. $\beta$ (a–d of Figure 1): As $\beta$ decreases (moving right), privacy increases and all methods lose test accuracy. For large $\beta$ (near the no-privacy case), all three algorithms attain accuracies close to the noiseless model. As $\beta$ tightens, the SR-PAC curve remains strictly above the Auto-PAC and Efficient-PAC curves across datasets. On CIFAR-10 and CIFAR-100, Auto-PAC and Efficient-PAC are visibly separated from each other (not merely from SR-PAC), reflecting their different Gaussian calibrations. On MNIST and AG-News, the three methods cluster near the top for larger $\beta$, but SR-PAC retains a measurable accuracy edge at matched $\beta$.
@@
Noise magnitude vs. $\beta$ (e–h of Figure 1): As $\beta$ decreases, each algorithm must add more noise, so all three curves rise. Across all datasets, SR-PAC uses the smallest $\mathbb{E}\bigl{[}\|B\|^{2}_{2}\bigr{]}$ at each $\beta$. Auto-PAC and Efficient-PAC both overshoot—they inject more noise than SR-PAC at matched $\beta$—and on CIFAR-100, MNIST and AG-News, they diverge from each other as well.